For a particular regression​ analysis, SSTequals=147.7251 and SSRequals=142.8402 a. Obtain the coefficient of determination. b. Determine SSE.
Answer #1

A) Given data

For a regression analysis

SST=147.7251

SSR=142.8402

Total sum of squares(SST) =Sum of squares regression (SSR) +Sum of squares of error(SSE)

a) Coefficient of determination (R2) is

R2=SSR/(SSR+SSE) =SSR/SST

=142.8402/(142.8402+4.8849)

=142.8402/147.7251 =0.966

b) Sum of squares of error (SSE)

SST=SSR+SSE

Now, SSE=SST-SSR

=147.7251-142.8402

=4.8849.